# F-95000 Cergy-Pontoise, France

### v1

@@ 2.2 The m-Chase Algorithm
Clearly to apply the $m\_Chase$-based approach described above, one has to answer the following questions:
  • Does the $m\_Chase$ algorithm terminate?
@@
  • Does the result contain all true tuples and all conflicting tuples that the dependencies can derive? In other words: which underlying semantics ensure that $m\_Chase$ algorithm is correct?
All these questions find positive answers in LauS25 , based on the set theoretic semantics introduced in CKS86 ; Spyratos87 , under the assumption that the set $FD$ is normalized. Following LauS25 , if $FD^{+}$ denotes the closure of $FD$ under the Armstrong’s axioms Armstrong , then $FD$ is said to be normalized if it contains all dependencies in $FD^{+}$ such that:
@@
In this section, we recall from LauS25 the basic formalism on which the algorithm $m\_Chase$ relies, namely that of multi-valued tuple. A multi-valued tuple, or m-tuple, extends the notion of tuple in the sense that an m-tuple associates every attribute $A$ with a possibly empty subset of the active domain of $A$ as opposed to a single value from the active domain.
@@
We call m-table over $U$ any finite set of m-tuples over $U$. For all $\sigma$ and $\sigma^{\prime}$ in an m-table $\Sigma$, and $X\to A$ such that $XA\subseteq sch(\sigma)$ and $XA\subseteq sch(\sigma^{\prime})$, the following rule called m-Chase rule generalizes the chase rule.
$\bullet$ m-Chase rule: Let $\sigma_{1}=\sigma\cup\sigma^{\prime}(A)$ and $\sigma^{\prime}_{1}=\sigma^{\prime}\cup\sigma(A)$
@@
As shown in Algorithm 1, our algorithm consists in applying the above m-Chase rule whenever ${\sf tuples}(\sigma(X))\cap{\sf tuples}(\sigma^{\prime}(X))\neq\emptyset$ until no further transformation is possible. The output is an m-table denoted by $m\_Chase(T)$. It has been shown in LauS25 that this algorithm always terminates and that the partition semantics of tuples in $\mathcal{T}$ (as introduced in Spyratos87 and extended in JIIS ; LauS25 ), can be defined based on $m\_Chase(T)$ as follows:
@@
0:  An m-table denoted by $m\_Chase(T)$.
1:  $\Sigma:=\{\sigma_{t}~{}|~{}t\in T\}$ // $\sigma_{t}$ is the m-tuple such that $\sigma_{t}(A)=\{t.A\}$ for $A\in sch(t)$
@@
8:              apply the m-Chase rule to $\sigma$ and $\sigma^{\prime}$
@@
10:  $m\_Chase(T):=\Sigma$
11:  return  $m\_Chase(T)$
@@
  1. 1.
    A tuple $t$ in $\mathcal{T}$ is in ${\sf True}(\mathcal{T})$ if and only if there exists $\sigma$ in $m\_Chase(T)$ such that:
    -- $sch(t)\subseteq sch(\sigma)$ (i.e., $\sigma$ has nonempty components over attributes on which $t$ is defined),
    -- $t\sqsubseteq\sigma$.
  2. 2.
    A tuple $t$ in $\mathcal{T}$ is in ${\sf Confl}(\mathcal{T})$ if and only if there exists $\sigma$ in $m\_Chase(T)$ such that:
    -- $sch(t)\subseteq sch(\sigma)$ and $t\sqsubseteq\sigma$,
    -- there exists $X\to A$ in $FD$ such that $XA\subseteq sch(t)$ and $|{\sf tuples}(\sigma(A)|>1$.
  3. 3.
    A tuple $t$ in $\mathcal{T}$ is in ${\sf Cons}(\mathcal{T})$ if and only if there exists $\sigma$ in $m\_Chase(T)$ such that:
    -- $sch(t)\subseteq sch(\sigma)$ and $t\sqsubseteq\sigma$,
    -- for every $X\to A$ in $FD$ such that $XA\subseteq sch(t)$, $|{\sf tuples}(\sigma(A))|=1$.
  4. 4.
    For every $\sigma$ in $m\_Chase(T)$ and every $S\subseteq sch(\sigma)$, either ${\sf tuples}(\sigma(S))\subseteq{\sf Cons}(\mathcal{T})$ or ${\sf tuples}(\sigma(S))\subseteq{\sf Confl}(\mathcal{T})$. $\Box$
As shown in LauS25 , the computation of $m\_Chase(T)$ is in $\mathcal{O}(|m\_Chase(T)|^{3}.\delta^{2})$, where $\delta$ is the maximal cardinality of the components of m-tuples in $m\_Chase(T)$, which is precisely the maximum number of $A$-values associated with $X$-values when $X\to A$ is a functional dependency in $FD$. As Algorithm 1 shows that $|m\_Chase(T)|\leq|T|$, we state that the computation of $m\_Chase(T)$ is in $\mathcal{O}(|T|^{3}.\delta^{2})$, i.e., polynomial in the size of $T$.
To illustrate Algorithm 1 and Proposition 1, consider again the context of Example 2 where $U=\{A,B,C\}$, $FD=\{A\to C,B\to C\}$ and $T=\{ab,bc,ac^{\prime}\}$. Running Algorithm 1 yields the following steps:
  • The algorithm starts with the m-table $\Sigma=\{(a)(b),(b)(c),(a)(c^{\prime})\}$
  • Applying $B\to C$ to the first two m-tuples, we obtain $\Sigma=\{(a)(b)(c),(a)(c^{\prime})\}$
@@
  • ${\sf Confl}(\mathcal{T})=\{abc,abc^{\prime},ac,ac^{\prime},bc,bc^{\prime}\}$.
@@
In the following section, we first define the notion of star schema and then we show that the results from LauS25 that have just been recalled apply in this context as well.
@@ 3.1 The Context of our Approach
We first recall from the literature Ullman that a star schema, as considered in our approach, consists of the following tables and constraints:
@@
  • a fact table $F$ defined over $K_{1},\ldots,K_{n},M_{1},\ldots,M_{p}$. The attributes $M_{1},\ldots,M_{p}$ are called measures, and we denote by $\mathbf{M}$ the set of all measures, that is $\mathbf{M}=\{M_{1},\ldots,M_{p}\}$.
@@
    In other words, for $i=1,\ldots,n$, $K_{i}$ is the key of $D_{i}$ and $K_{1}\ldots K_{n}$ is the key of $F$. We denote by $\mathbf{K}$ the set of all dimension keys that is $\mathbf{K}=\{K_{1},\ldots,K_{n}\}$.
It is easy to see that if $FD$ is defined as above, then for every non trivial functional dependency $X\to A$ in $FD^{+}$ we have: $X\cap\mathbf{K}\neq\emptyset$. More precisely, if $A$ is in $\mathbf{K}$ then $A$ must occur in $X$, in which case $X\to A$ is trivial (because $FD$ contains no dependency whose right hand-side is in $\mathbf{K}$), if $A$ is in $sch^{*}(D_{i})$ then $X$ must contain $K_{i}$ and if $A$ is in $\mathbf{M}$ then $X$ must contain $\mathbf{K}$. Thus, for every non trivial functional dependency $X\to A$ in $FD^{+}$, there exists $X_{0}\to A$ in $FD$ such that $X_{0}\subseteq X$.
Since the left hand-sides of the dependencies in $FD$ can not be reduced further, this means that $FD$ is normalized. On the other hand as the left hand-sides of functional dependencies in $FD$ are attributes in $\mathbf{K}$, not occurring in the right hand-sides of these dependencies, $FD$ is acyclic. As a consequence, all results in LauS25 apply in the context of star schemas.
@@
    For every $t$ in $F$, $t$ is defined over $\mathbf{K}$ and over at least one measure attribute in $\mathbf{M}$. We consider that storing a fact with no associated measure value makes no sense.
Example 3
We illustrate the above concepts using a toy example that will serve as a running example in the remainder of this paper. We consider two dimensions $D_{1}$ and $D_{2}$ such that $sch(D_{1})=K_{1}A_{1}^{1}A_{1}^{2}$ and $sch(D_{2})=K_{1}A_{2}^{1}A_{2}^{2}$. Moreover the fact table $F$ is such that $sch(F)=K_{1}K_{2}M_{1}$ meaning that we consider one measure attribute $M_{1}$. As specified above, we have $FD=\{K_{1}\to A_{1}^{1},K_{1}\to A_{1}^{2},K_{2}\to A_{2}^{1},K_{2}\to A_{2}^{2},K_{1}K_{2}\to M_{1}\}$. The content of the tables $D_{1}$, $D_{2}$ and $F$ is shown in the following Figure 1.
@@
We also stress that the key value $k^{\prime\prime}_{1}$ occurs in $D_{1}$ but not in $F$, whereas the key value $k^{\prime\prime}_{2}$ over $K_{2}$ occurs in $F$ but not in $D_{2}$. These two cases respectively illustrate that key values in a dimension table may not occur in the fact table and that the foreign key constraint between a dimension table and the fact table may not be satisfied (contrary to what is generally assumed in the literature). $\Box$
3.2 True, Consistent and Conflicting Tuples in a Star Schema
The following proposition states important properties regarding m-tuples of the table $m\_Chase(T)$ where $T$ is the table collecting all tuples in the data warehouse. In the remainder of this paper, we refer to such table as a star-table.
Proposition 2
Let $T$ be a star-table over universe $U$. The following hold:
  1. 1.
    For every $\sigma$ in $m\_Chase(T)$ and every $i=1,\ldots,n$, if $K_{i}\in sch(\sigma)$ then $|{\sf tuples}(\sigma(K_{i}))|=1$. Consequently, if $\mathbf{K}\subseteq sch(\sigma)$ then $|{\sf tuples}(\sigma(\mathbf{K}))|=1$.
  2. 2.
    For every tuple $k$ over $\mathbf{K}$ in $\mathcal{T}$, there exists at most one $\sigma$ in $m\_Chase(T)$ such that $\mathbf{K}\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(\mathbf{K}))=\{k\}$.
  3. 3.
    Moreover, $m\_Chase(T)$ contains the following two kinds of m-tuples:
    1. (a)
      $\sigma$ for which there exists $i_{0}\in\{1,\ldots,n\}$ such that:
      -- $sch(\sigma)\subseteq sch(D_{i_{0}})$, ${\sf tuples}(\sigma(K_{i_{0}}))=\{k_{i_{0}}\}$ and for every $t\in F$, $t.K_{i_{0}}\neq k_{i_{0}}$.
      -- for every $A\in sch^{*}(D_{i_{0}})$, $\sigma(A)=\{a~{}|~{}(\exists q\in D_{i})(q.K_{i_{0}}=k_{i_{0}}\wedge q.A=a)\}$
    2. (b)
      $\sigma$ such that $\mathbf{K}\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(\mathbf{K}))=\{k\}$, and
      -- for every $M_{j}\in\mathbf{M}$, $\sigma(M_{j})=\{m_{j}~{}|~{}(\exists t\in F)(t.\mathbf{K}=\sigma(\mathbf{K})\wedge t.M_{j}=m_{j})\}$
      -- for every $i=1,\ldots,n$, for every $A\in sch^{*}(D_{i})$, $\sigma(A)=\{a~{}|~{}(\exists t\in D_{i})(t.K_{i}=k.K_{i}\wedge t.A=a)\}$.
Proof. 1. This result comes from the fact that, in order to generate multi-valued components of an attribute $A$, the algorithm $m\_Chase$ has to consider a functional dependency whose right hand-side is $A$. As in the case of a star-table, no dependency in $FD$ has its right hand-side in $\mathbf{K}$, the proof of this item is complete.
2. Let $\sigma_{1}$ and $\sigma_{2}$ be in $m\_Chase(T)$ such that $\mathbf{K}\subseteq sch(\sigma_{i})$ for $i=1,2$ and $\sigma_{1}(\mathbf{K})=\sigma_{2}(\mathbf{K})$. Since $\sigma_{1}\neq\sigma_{2}$ (the algorithm eliminates duplicates) there exists $A$ not in $\mathbf{K}$ such that $\sigma_{1}(A)\neq\sigma_{2}(A)$. Since $A$ is not in $\mathbf{K}$, either there is $i_{0}$ in $\{1,\ldots,n\}$ such that $A\in sch^{*}(D_{i_{0}})$ or $A$ is in $\mathbf{M}$. In either case, $FD$ contains a dependency of the form $X\to A$ where $X=K_{i_{0}}$ in the former case, or $X=\mathbf{K}$ in the latter case. Since we have $\sigma_{1}(X)=\sigma_{2}(X)$, applying $m\_Chase$ to $m\_Chase(T)$ changes $\sigma_{1}(A)$ and $\sigma_{2}(A)$ into $\sigma_{1}(A)\cup\sigma_{2}(A)$. As by definition of $m\_Chase(T)$, we have $m\_Chase(m\_Chase(T))=m\_Chase(T)$, we obtain a contradiction which completes the proof.
3. We first notice that, as stated above, for every $K_{i}$ in $\mathbf{K}$ and every $\sigma$ in $m\_Chase(T)$ such that $K_{i}$ is in $sch(\sigma)$, we have $|{\sf tuples}(\sigma(K_{i}))|=1$. This is so because no functional dependency in $FD$ has a right hand-side in $\mathbf{K}$, which makes it impossible to generate conflicts on these attributes.
Moreover, the m-tuples of the first kind are the result of joining two rows $\rho_{1}$ and $\rho_{2}$ in the current table such that $sch(\rho_{1})=sch(\rho_{2})=sch(D_{i_{0}})$, $\rho_{1}(K_{i_{0}})=\rho_{2}(K_{i_{0}})$, and this key value does not occur in $F$ thus preventing these m-tuples to be combined with m-tuples over $sch(F)$. Similarly, the tuples of the second kind are obtained by joining every tuple $km$ in $F$ with tuples $t_{i}$ in $D_{i}$ such that $k.K_{i}=t_{i}.K_{i}$ ($i=1,\ldots,n$). We thus obtain that the set $\Sigma$ of all m-tuples as described above occur in $m\_Chase(T)$.
To end the proof we now argue that applying the $m\_Chase$ procedure to $\Sigma$ has no effect. This is so because for every $X\to A$ in $FD$ and all rows $\rho_{1}$ and $\rho_{2}$ in $\Sigma$ such that $\rho_{1}(X)\cap\rho_{2}(X)\neq\emptyset$, the $X$-values of these rows are reduced to one tuple and thus we have $\rho_{1}(X)=\rho_{2}(X)$. By definition of $\Sigma$, it follows that $\rho_{1}(A)=\rho_{2}(A)$, which completes the proof. $\Box$

### v2

@@ 2.3 The m-Chase Algorithm
Clearly to apply the m_Chasem\_Chase-based approach described above, one has to answer the following questions:
  • Does the m_Chasem\_Chase algorithm terminate?
@@
  • Does the result contain all true tuples and all conflicting tuples that the dependencies can derive? In other words: which underlying semantics ensure that m_Chasem\_Chase algorithm is correct?
All these questions find positive answers in LauS25 , based on the set theoretic semantics introduced in CKS86 ; Spyratos87 , under the assumption that the set $FD$ is normalized. Following LauS25 , given a set of attributes $U$ and a set of functional dependencies $FD$ over $U$, if $FD^{+}$ denotes the closure of $FD$ under the Armstrong’s axioms Armstrong , then $FD$ is said to be normalized if it contains all dependencies in $FD^{+}$ such that:
@@
In this section, we recall from LauS25 the basic formalism on which the algorithm m_Chasem\_Chase relies, namely that of multi-valued tuple. A multi-valued tuple, or m-tuple, extends the notion of tuple in the sense that an m-tuple associates every attribute $A$ with a possibly empty subset of the active domain of $A$ as opposed to a single value from the active domain.
@@
Similarly, the set theoretic union operation is extended to m-tuples in a ‘component-wise’ manner, that is: the union of m-tuples $\sigma$ and $\sigma^{\prime}$ is the m-tuple denoted by $\sigma\sqcup\sigma^{\prime}$ and for every $A$ in $U$, $(\sigma\sqcup\sigma^{\prime})(A)=\sigma(A)\cup\sigma^{\prime}(A)$.
For example, if $A$, $B$ and $C$ are three attributes, $\sigma=(aa^{\prime})(b)$ and $\sigma^{\prime}=(a)$ are two m-tuples such that $sch(\sigma)=AB$ and $sch(\sigma^{\prime})=A$. In this case, we have $\sigma^{\prime}\sqsubseteq\sigma$ because $\sigma^{\prime}(A)\subseteq\sigma(A)$ and $\sigma^{\prime}(B)=\emptyset$ whereas $\sigma(B)=(b)$, and so, it holds that $\sigma\sqcup\sigma^{\prime}=\sigma$. Moreover, if $\sigma^{\prime\prime}=(a)(bb^{\prime})(c)$, then $sch(\sigma^{\prime\prime})=ABC$ and we have $\sigma\sqcup\sigma^{\prime\prime}=(aa^{\prime})(bb^{\prime})(c)$.
We call m-table over $U$ any finite set of m-tuples over $U$. For all $\sigma$ and $\sigma^{\prime}$ in an m-table $\Sigma$, and for every functional dependency $X\to A$ over $U$ such that $XA\subseteq sch(\sigma)$ and $XA\subseteq sch(\sigma^{\prime})$, the following rule called m_Chase rule generalizes the chase rule.
$\bullet$ m_Chase rule: Let $\sigma_{1}=\sigma\sqcup\sigma^{\prime}(A)$ and $\sigma^{\prime}_{1}=\sigma^{\prime}\sqcup\sigma(A)$
@@
As shown in Algorithm 1, our approach consists in applying the above m_Chase rule whenever ${\sf tuples}(\sigma(X))\cap{\sf tuples}(\sigma^{\prime}(X))\neq\emptyset$ until no further transformation is possible. According to this algorithm, at each step, if $\sigma$ and $\sigma^{\prime}$ are two m-tuples in the current state of $\Sigma$ whose schemas contain $XA$ where $X\to A$ is a functional dependency in $FD$, if their $X$-value share at least one tuple then their $A$-values are set to the same value $\sigma(A)\cup\sigma^{\prime}(A)$ (see the first and last line of the above m_Chase rule), and in case these changes imply that one m-tuple subsumes the other one, the redundancy is removed (as stated in the two intermediate lines of the above m_Chase rule). Similarly to the $Chase$ algorithm, this process is applied in Algorithm 1, until no further transformation of $\Sigma$ is possible. The output of Algorithm 1 is then an m-table denoted by m_Chase(T)m\_Chase(T).
To illustrate this step in the context of Example 2, consider first $\sigma=(a)(b)$ and $\sigma^{\prime}=(b)(c)$ with the dependency $B\to C$. In this case $\sigma_{1}=(a)(b)(c)$ and $\sigma^{\prime}_{1}=\sigma^{\prime}$. As $\sigma^{\prime}_{1}\sqsubseteq\sigma_{1}$, only $\sigma_{1}$ is kept for the next computation step. Thus, for $\sigma=(a)(b)(c)$ and $\sigma^{\prime}=(a)(c^{\prime})$ with the dependency $A\to C$, we have $\sigma_{1}=(a)(b)(cc^{\prime})$ and $\sigma^{\prime}_{1}=(b)(cc^{\prime})$. In this case again, we have $\sigma^{\prime}_{1}\sqsubseteq\sigma_{1}$, and so, $\sigma^{\prime}_{1}$ is discarded, and we are left with the only m-tuple $(a)(b)(cc^{\prime})$. Since no further application of the m_Chase rule is possible, we obtain that $m\_Chase(T)=\{(a)(b)(cc^{\prime})\}$.
[…]
It has been shown in LauS25 that Algorithm 1 always terminates and that the partition semantics of tuples in $\mathcal{T}$ (as introduced in Spyratos87 and extended in JIIS ; LauS25 ), can be defined based on m_Chase(T)m\_Chase(T) according to the following proposition. To state these basic results, we use the following notation given a table $T$: the sets of all true tuples in $\mathcal{T}$, all conflicting tuples in $\mathcal{T}$, all consistent tuples in $\mathcal{T}$ are respectively denoted by ${\sf True}(\mathcal{T})$, ${\sf Confl}(\mathcal{T})$ and ${\sf Cons}(\mathcal{T})$.
@@
0: An m-table denoted by m_Chase(T)m\_Chase(T).
[…]
@@
8:     apply the m_Chase rule to $\sigma$ and $\sigma^{\prime}$
@@
[…]
11:return m_Chase(T)m\_Chase(T)
@@
  • A tuple $t$ in $\mathcal{T}$ is in ${\sf True}(\mathcal{T})$ if and only if there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that:
    - $sch(t)\subseteq sch(\sigma)$ (i.e., $\sigma$ has nonempty components over attributes in $sch(t)$),
    - $t\sqsubseteq\sigma$.
  • A tuple $t$ in $\mathcal{T}$ is in ${\sf Confl}(\mathcal{T})$ if and only if there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that:
    - $sch(t)\subseteq sch(\sigma)$ and $t\sqsubseteq\sigma$,
    - there exists $X\to A$ in $FD$ such that $XA\subseteq sch(t)$ and $|{\sf tuples}(\sigma(A)|>1$.
  • A tuple $t$ in $\mathcal{T}$ is in ${\sf Cons}(\mathcal{T})$ if and only if there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that:
    - $sch(t)\subseteq sch(\sigma)$ and $t\sqsubseteq\sigma$,
    - for every $X\to A$ in $FD$ such that $XA\subseteq sch(t)$, $|{\sf tuples}(\sigma(A))|=1$.
  • For every $\sigma$ in m_Chase(T)m\_Chase(T) and every $S\subseteq sch(\sigma)$, either ${\sf tuples}(\sigma(S))\subseteq{\sf Cons}(\mathcal{T})$ or ${\sf tuples}(\sigma(S))\subseteq{\sf Confl}(\mathcal{T})$. $\Box$
We notice that when ${\sf Confl}({T})=\emptyset$, i.e., when no tuples in $\mathcal{T}$ are conflicting, then $T$ satisfies all functional dependencies in $FD$, which is denoted by $T\models FD$. In this case, as shown by Proposition 1, all components of all m-tuples in m_Chase(T)m\_Chase(T) contain one value, and so, m_Chase(T)m\_Chase(T) can be seen as a table, which is in fact the table $Chase(T)$.
Regarding the time complexity of Algorithm 1, it has been shown in LauS25 that the computation of m_Chase(T)m\_Chase(T) is in $\mathcal{O}(|m\_Chase(T)|^{3}\cdot\delta^{2})$, where $\delta$ is the maximal cardinality of the components of m-tuples in m_Chase(T)m\_Chase(T), which is precisely the maximum number of $A$-values associated with $X$-values when $X\to A$ is a functional dependency in $FD$. As Algorithm 1 shows that $|m\_Chase(T)|\leq|T|$, we state that the computation of m_Chase(T)m\_Chase(T) is in $\mathcal{O}(|T|^{3}\cdot\delta^{2})$, i.e., polynomial in the size of $T$. We notice that in this complexity result, the size of the universe $U$ is a constant, contrary to the size of m-tuples. In fact the size of every m-tuple in m_Chase(T)m\_Chase(T) is bounded by $\delta\cdot|U|$. As will be seen shortly, these bounds can be refined when considering star-schemas.
[…]
Thus, using Proposition 1, the set ${\sf Confl}_{\min}(\mathcal{T})$ is characterized as follows: $t$ is in ${\sf Confl}_{\min}(\mathcal{T})$ if and only if one of the following two statements holds:
  • there exist $\sigma$ in m_Chase(T)m\_Chase(T), $i_{0}$ in $\{1,\ldots,n\}$ and $A$ in $sch^{*}(D_{i_{0}})\cap sch(\sigma)\cap sch(t)$ such that $t=k_{i_{0}}a$ is in ${\sf tuples}(\sigma(K_{i_{0}}A))$ and $|{\sf tuples}(\sigma(A))|>1$,
  • there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that $\mathbb{K}\subseteq sch(\sigma)\cap sch(t)$, there exists $M_{i}$ in $\mathbb{M}\cap sch(\sigma)\cap sch(t)$ such that $t=km_{i}$ is in ${\sf tuples}(\sigma(\mathbb{K}M_{i}))$ and $|{\sf tuples}(\sigma(M_{i}))|>1$.
To illustrate Algorithm 1 and Proposition 1, consider again the context of Example 2 where $U=\{A,B,C\}$, $FD=\{A\to C,B\to C\}$ and $T=\{ab,bc,ac^{\prime}\}$. As seen earlier, running Algorithm 1 yields the following steps:
- The algorithm starts with the m-table $\Sigma=\{(a)(b),(b)(c),(a)(c^{\prime})\}$.
- Applying $B\to C$ to the first two m-tuples, we obtain $\Sigma=\{(a)(b)(c),(a)(c^{\prime})\}$.
@@
  • ${\sf Confl}_{min}(\mathcal{T})=\{ac,ac^{\prime},bc,bc^{\prime}\}$ and so, we have ${\sf Confl}(\mathcal{T})=\{abc,abc^{\prime},ac,ac^{\prime},bc,bc^{\prime}\}$.
@@
In the following section, we first recall the notion of star schema and then we show that the results from LauS25 that have just been recalled apply in this context as well.
@@ 3.1 The Context of our Approach
We first recall from the literature inmon96 ; Ullman that a star schema, as considered in our approach, consists of the following tables and constraints:
@@
  • a fact table $F$ defined over $K_{1},\ldots,K_{n},M_{1},\ldots,M_{p}$. The attributes $M_{1},\ldots,M_{p}$ are called measures, and we denote by $\mathbb{M}$ the set of all measures, that is $\mathbb{M}=\{M_{1},\ldots,M_{p}\}$.
@@
    In other words, for $i=1,\ldots,n$, $K_{i}$ is the key of $D_{i}$ and $K_{1}\ldots K_{n}$ is the key of $F$. We denote by $\mathbb{K}$ the set of all dimension keys that is $\mathbb{K}=\{K_{1},\ldots,K_{n}\}$.
It is easy to see that if $FD$ is defined as above, then for every non trivial functional dependency $X\to A$ in $FD^{+}$, it holds that $X\cap\mathbb{K}\neq\emptyset$ and $A\not\in\mathbb{K}$. More precisely, given an attribute $A$, only one of the following three cases is possible: $(i)$ AA is in $\mathbb{K}$, or $(ii)$ AA is in $sch^{*}(D_{i})$ for some $i=1,\ldots,n$, or $(iii)$ AA is in $\mathbb{M}$. For each of these cases, we have the following: $(i)$ if $A$ is in $\mathbb{K}$ then $A$ must occur in $X$, in which case $X\to A$ is trivial (because $FD$ contains no dependency whose right hand-side is in $\mathbb{K}$); $(ii)$ if $A$ is in $sch^{*}(D_{i})$ then $X$ must contain $K_{i}$; $(iii)$ if $A$ is in $\mathbb{M}$ then $X$ must contain $\mathbb{K}$. Thus, for every non trivial functional dependency $X\to A$ in $FD^{+}$, there exists $X_{0}\to A$ in $FD$ such that $X_{0}\subseteq X$.
Since the left hand-sides of the dependencies in $FD$ can not be reduced further, this means that $FD$ is normalized. On the other hand as the left hand-sides of functional dependencies in $FD$ are attributes in $\mathbb{K}$, not occurring in the right hand-sides of these dependencies, $FD$ is acyclic. As a consequence, all results in LauS25 apply in the context of star schemas.
@@
    For every $t$ in $F$, $t$ is defined over $\mathbb{K}$ and over at least one measure attribute in $\mathbb{M}$. We consider that storing a fact in which some key values are missing makes no sense, since we consider that measure values depend on all dimensions. Similarly, considering a fact with no associated measure value makes no sense.
We notice that the above restrictions follow common sense considerations (storing values without corresponding key values or storing key values alone with no other attribute values make no sense). However, we also notice that relaxing these restrictions can be supported by our approach, at the cost of technical complications.
@@
Example 3
We illustrate the above concepts using a toy example that will serve as a running example in the remainder of this paper. We consider two dimensions $D_{1}$ and $D_{2}$ such that $sch(D_{1})=K_{1}A_{1}^{1}A_{1}^{2}$ and $sch(D_{2})=K_{2}A_{2}^{1}A_{2}^{2}$. Moreover the fact table $F$ is such that $sch(F)=K_{1}K_{2}M_{1}$ meaning that we consider one measure attribute $M_{1}$. As specified above, we have $FD=\{K_{1}\to A_{1}^{1},K_{1}\to A_{1}^{2},K_{2}\to A_{2}^{1},K_{2}\to A_{2}^{2},K_{1}K_{2}\to M_{1}\}$. The content of the tables $D_{1}$, $D_{2}$ and $F$ is shown in Figure 1.
@@
We also stress that the above two restrictions on missing values are satisfied and that key value $k^{\prime\prime}_{1}$ occurs in $D_{1}$ but not in $F$, whereas the key value $k^{\prime\prime}_{2}$ over $K_{2}$ occurs in $F$ but not in $D_{2}$. These two cases respectively illustrate that key values in a dimension table may not occur in the fact table and that the foreign key constraint between a dimension table and the fact table may not be satisfied (contrary to what is generally assumed in the literature).